1. There are three workers in a company (Van, Bondy, Larry)

-in one hour Van can mow 1 lawn or wash 1 car

-in one hour Bondy can mow 1 lawn or wash 2 cars

-in one hour Larry can mow 2 lawns or wash 1 car

In a 5 hour day the three of them together can mow a total of ____ lawns.

In a 5 hour day the three of them together can wash a total of ____ cars.

Suppose the three all devote 4 hours to mowing lawns and 1 hour to washing cars. In this case they can mow ____ lawns and wash _____ cars.

2. A production possibilities frontier identifies combinations of two goods that can be produced with given resources and technology. T/F
Answer #1

1) In 5 hours,Van can mow (5*1)=5 lawns

In 5 hours, Bondy can mow (5*1)=5 lawns

In 5 hours,Lary can mow (5*2)=10 lawns

So in 5 hours, all of them can together mow 20 lawns.

In 5 hours,Van can wash (5*1)=5 cars

In 5 hours, Bondy can wash (5*2)=10 cars

In 5 hours,Lary can wash (5*1)=5 cars

So in 5 hours, all of them can together wash 20 cars.

In 4 hours,Van can mow (4*1)=4 lawns

In 4 hours, Bondy can mow (4*1)=4 lawns

In 4 hours,Lary can mow (4*2)=8 lawns

So in 4 hours, all of them can together mow 16 lawns.

In 1 hours Van can wash 1 car , Bondy can wash 2 cars and Lary can wash 1 car

So all three can together wash 4 cars.

So, with 4 hours of mowing and 1 hour of washing cars, these three can together mow 16 lawns and wash 4 cars.

2)It is true that a production possibility frontier identifies combinations of two goods thatcan ve produced with giev resources and technology. It connects all the combinations of two goods that an economy can produce efficiently. So, answer is true.